OpenStudy (anonymous):

solve the rational expression 2x+5 2x 1 ----- - ----- = - --- 6 3 2

OpenStudy (anonymous):

\[\frac{ 2x + 5 - 4x }{ 6 }=\frac{ -3 }{ 6}\] ==> \[\frac{ -2x+5 }{ 6 }=\frac{ -3 }{ 6 }\] ==> -2x=-3-5 -2x=-8 x=4 give medal plz